Exercise 2-13 Departmental Predetermined Overhead Rates [LO2-1, LO2-2, L02-4] White Company has two departments, Cutting and Finishing. The company uses a job-order costing system and computes a predetermined overhead rate in each department. The Cutting Department bases its rate on machine-hours, and the Finishing Department bases its rate on direct labor-hours. At the beginning of the year, the company made the following estimates Department Cutting Finishing Direct labor-hours Machine-hours Total fixed manufacturing overhead cost Variable manufacturing overhead per machine-hour Variable manufacturing overhead per direct labor-hour 78,000 2,100 $ 380,000 $ 504,000 7,100 58,500 $ 3.00 $ 4.75 Required: 1. Compute the predetermined overhead rate for each department. 2. The job cost sheet for Job 203, which was started and completed during the year, showed the following: Department Cutting Finishing Direct labor-hours Machine-hours Direct materials Direct labor cost 12 83 $790 $115 $370 $276 Using the predetermined overhead rates that you computed in requirement (1), compute the total manufacturing cost assigned to Job 203 3. Would you expect substantially different amounts of overhead cost to be assigned to some jobs if the company used a plantwide predetermined overhead rate based on direct labor-hours, rather than using departmental rates? Complete the question by entering your answers in the tabs given below. Required 1 Required 2 Required 3 Compute the predetermined overhead rate for each department. (Round your answers to 2 decimal places.) Department Cutting Finishing Predetermined overhead rate per MH r DLHRequired 1Required 2 Required 3 Using the predetermined overhead rates that you computed in requirement (1), compute the total manufacturing cost assigned to Job 203. (Round intermediate calculations to 2 decimal places and final answer to the nearest dollar amount. Total manufacturing cost K Required 1 Required 3
Answer #1

1) Predetermine overhead rate :

Cutting Finishing
Predetermine overhead rate (380000/58500+3) = $9.50 per machine hour (504000/78000+4.75) = $11.21 per labor hour

2) Calculate total manufacturing cost

Cutting Finishing
direct material 790 370
Direct labor 115 276
Overhead applied 83*9.50 = 788.50 12*11.21 = 134.52
1694 781

Total manufacturing cost = 1694+781 = $2475